A regid body sping with angular velocity of 4 radias per sec about axis parallel to (3j-k) passing through the point (I+3j+k) .find find the velocity of the partical at a point (4i-2j+k)
 

Grade:12th pass

1 Answers

Arun
4 years ago
 

Obviously,the unit vector (^n) in the direction of

                              3j – k = (3j – k )/l 3j – k l

                                                         =(3j – k)/sqrt(10)

angular velocity of a rigid body= 4*[(3j – k )/sqrt(10)]= 12j – 4k / sqrt(10)

Let the point whose velocity is desired to be determined be P.Then its +ve w.r.t. the point 2i-j-k on the axis given by

r=(4i – 2j + k)-(i + 3j + k)= 3i – 5j

Hence,if v be the velocity of pt. P,we have v=w(omegha)cross r

  = (3j – k) cross (3i – 5j) =  – 5i – 3j– 9k